OpenStudy (anonymous):

how do I use the half-angle formula to determine the exact value of the given trigonometric expression? cot(3pi/8)

OpenStudy (anonymous):

Assumming you only have half-angle formula for sin and cos. What is cot(x) in terms of cos(x) and sin(x)?

OpenStudy (anonymous):

cosx/sinx

OpenStudy (anonymous):

What are the half angle formulae for those sin and cos?

OpenStudy (anonymous):

\[\cos(x/2)=\sqrt{(1+cosx)/2} , \sin(x/2)= \sqrt{(1-cosx)/2} \]

OpenStudy (raden):

the half-angle formula for sin, cos, and tan u can look here : http://www.freemathhelp.com/trig-double-angles.html also, (3pi/8) if u converting to degrees it can be 135/2 degrees